OpenStudy (anonymous):

What is the reciprocal of 4? A.-1/4 B.1 C.4 D.1/4

OpenStudy (anonymous):

brb

OpenStudy (astrophysics):

Reciprocal is just \[\frac{ 1 }{ original~number }\]

OpenStudy (haseeb96):

the reciprocal of a number suppose x is 1/x now what will be its answer

OpenStudy (phi):

when you multiply a number by its reciprocal you get 1 for fractions, you "flip" them for example , the reciprocal of 2/3 is 3/2 if you have a whole number such as 4, write it as 4/1 now you can use the "flip" rule... you get 1/4 notice \( \frac{1}{4} \cdot 4 = \frac{4}{4} = 1 \) showing that 4 and 1/4 are reciprocals (because when you multiply them, you get 1)